OpenStudy (anonymous):

Hi can anyone assist me factor this x^3 -9x^2

OpenStudy (anonymous):

what do u wanna do with this?

OpenStudy (anonymous):

x^3-9x^2=0 ???

OpenStudy (anonymous):

basically, is just to factorize it

OpenStudy (anonymous):

what do you think is the common term

OpenStudy (anonymous):

has no equals at the end, i find it confusing

OpenStudy (anonymous):

in x^3 and 9x^2 what can you take out

OpenStudy (anonymous):

no equals implies that it is just an expression when you have an equals sign it is an equation

OpenStudy (anonymous):

I am thinking 3 is common to both terms yes it's an equation

OpenStudy (anonymous):

if you have trouble seeing it like this try seeing it like this x* x* x - 9 x* x

OpenStudy (anonymous):

so on the left is x^3 and on the right is x^2

OpenStudy (anonymous):

yeah right

OpenStudy (anonymous):

what is common ?

OpenStudy (anonymous):

x is common

OpenStudy (anonymous):

bingo next step how many x's are common

OpenStudy (anonymous):

2 pairs of x?

OpenStudy (anonymous):

YES!!!

OpenStudy (anonymous):

so i will just take the two pairs of x out x*x (x-9) => x*x is just x^2 x^2(x-9)

OpenStudy (anonymous):

do you get what i did this is very basic stuff, factoring gets harder so let me know if you didnt get it

OpenStudy (anonymous):

i'm sorry that confused me a bit

OpenStudy (anonymous):

its ok dw about it

OpenStudy (anonymous):

x*x*x=> x^3 9*x*x=>9x^2

OpenStudy (anonymous):

following me?

OpenStudy (anonymous):

yes i get that

OpenStudy (anonymous):

if you kinda line up the x s you will see that two of the xs are common right?

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

x^3=x^2(x) 9x^2=x^2(9) back to our equation we plugin the above x^3-9x^2 becomes x^2(x)-x^2(9)

OpenStudy (anonymous):

since x^2 is common we factor it out and it becomes x^2(x-9)

OpenStudy (anonymous):

okay i got that on my worksheet, just didnt know how i got it

OpenStudy (anonymous):

??

OpenStudy (anonymous):

you can not guess answers in math you must be sure of what you are doing

OpenStudy (anonymous):

explain to me what you are doing and i will correct you

OpenStudy (anonymous):

just the working was different

OpenStudy (anonymous):

do explain

OpenStudy (anonymous):

dont worry everywhere is here to learn =)

OpenStudy (anonymous):

im not judging you i just wanna make sure you are doing it right

OpenStudy (anonymous):

anyway i have stuff of my own to do essentially for factoring you wanna find what is common between the terms and take it out =) goodluck